# SMARTe-VR: Student Monitoring and Adaptive Response Technology for e-learning in Virtual Reality

### v1

@@ 4.3 Results
Figure 2 presents the results obtained for the proposed models compared with traditional IRT across various time windows ${W_{l}\in\{1,\ldots,8\}}$ minutes.
@@
Results indicated that with short time windows, traditional IRT outperformed the two deep learning-based models, which struggled in this range. However, from 4 minutes onward, facial features began to provide valuable insights, enhancing both models’ ability to predict student state and understanding beyond the capabilities of traditional IRT. Both models showed improved performance as $W_{l}$ increased, suggesting that facial features more accurately represented the student’s state over longer periods. With time windows of several minutes, the models proposed were better able to capture behavioral patterns in facial expressions. Additionally, given that Japanese students tend to maintain more neutral and consistent facial expressions compared to other cultures, facial changes were likely subtler and less frequent. This suggests that, with shorter time windows, the collected facial signals may not have sufficiently reflected the student’s comprehension state. SMART-MLP, based on global features, achieved the highest accuracy of 85% in the 8-minute window. Both models demonstrated similar performance across all time windows, though SMART-MLP performed slightly worse in shorter windows, likely due to the limitations of global features in capturing rapid patterns.
@@
Table 4: Accuracy obtained for the SMART-MLP model with a $W_{l}=8$ minutes and for the traditional IRT model, evaluated across different content difficulty levels labeled in the SMARTe-VR dataset.
Finally, for the SMART-MLP model in the 8-minute time window, which achieved the best results, we analyzed its accuracy by the content difficulty levels provided in the dataset (see Table 4). The SMART-MLP model outperformed the traditional IRT model across all content. Both models showed a similar trend, being more effective in detecting understanding for medium-level content, followed by difficult-level content, with the lowest accuracy for easy-level content. With more medium-difficulty sessions in the dataset, SMART-MLP can more easily identify patterns. Additionally, IRT models benefit from medium-difficulty items, as the variability in responses allows the model to better discriminate between skill levels, crucial for accurate estimates.
5 Conclusion and Future Work
@@
In this work, we have: i) reviewed adaptive learning models and datasets; ii) introduced a VR platform for immersive conference visualization in e-learning, specifically designed for adaptive learning, which monitors both biometric data and learning metadata; iii) released the 5 public VR-based educational dataset; iv) proposed five research challenges tied to this dataset; and v) designed two architectures for understanding detection to support dataset validation.
@@
We believe VR will transform education. SMARTe-VR provides a comprehensive VR solution equipped with tools such as Auto QA, textbook creation, and biometric monitoring, enhancing educator capabilities and enabling AI-driven advancements. The SMARTe-VR dataset and its associated challenges offer a valuable resource for advancing AL models in VR, supporting research in IRT, KT, and MMLA
@@
SMART models offer a novel approach based on state-of-the-art IRT models, with a focus on understanding detection during learning. By leveraging facial biometrics, they enhance adaptability, enable real-time adjustments based on student states, and reduce reliance on response data. These methods validated our dataset with 85% accuracy, although further research is needed to ensure broader applicability.
@@
In future work, we plan to expand the dataset with more users, additional questions per video, a broader range of lecture videos, and further sensor integrations to create a more robust resource. We also plan to monitor students in e-learning settings for direct comparisons.
@@
Furthermore, MR will be integrated into the platform to enhance the user experience. To address the challenges, we will share models with the community and refine SMART-MLP and SMART-TCN models to detect understanding and include attention and fatigue detection. Ultimately, our goal is to incorporate these models into SMARTe-VR, creating a comprehensive adaptive learning experience that leverages the potential of VR-based education.
6 Acknowledgments
@@
This study has been supported by project BBforTAI (PID2021-127641OB-I00 MICINN/FEDER). A. Morales is also supported by the Madrid Government under the Multiannual Agreement with UAM in the line of Excellence for the University Teaching Staff (V PRICIT). This study is particularly grateful to the experimenters from Nagoya University and Hiroshima City University for their significant contributions to data collection and experiment implementation. Additionally, we extend our thanks to both universities for providing experimental equipment and technical support, which were indispensable for the successful completion of this research. Without their support, this study would not have been possible.

### v3

@@ 4.3. Results
Fig. 3 presents the results obtained for the proposed models compared to traditional IRT for various time windows $W_{l}\in\{1,\ldots,8\}$} (in minutes).
@@
The results indicated that with short time windows, the traditional IRT outperformed the two deep learning-based models, which struggled in this range. However, from 4 minutes onward, facial features began to provide valuable insights, enhancing both models’ ability to predict student state and understanding beyond the capabilities of traditional IRT. Both models showed better performance as $W_{l}$ increased, suggesting that facial features more accurately represented the student’s state over longer periods. With time windows of several minutes, the proposed models were better able to capture behavioral patterns in facial expressions. Furthermore, given that Japanese students tend to maintain more neutral and consistent facial expressions compared to other cultures, facial changes were likely subtler and less frequent. This suggests that, with shorter time windows, the collected facial signals may not have adequately reflected the student’s comprehension state. SMART-MLP, based on global features, achieved the highest accuracy of 85% in the 8-minute window. Both models demonstrated similar performance across all time windows, though SMART-MLP performed slightly worse in shorter windows, likely due to the limitations of global features in capturing rapid patterns.
Finally, for the SMART-MLP model in the 8-minute time window, which achieved the best results, we analyzed its accuracy by the content difficulty levels provided in the dataset (see Table 3). The SMART-MLP model outperformed the traditional IRT model in all content. Both models showed a similar trend, being more effective in detecting understanding for medium-level content, followed by difficult-level content, with the lowest accuracy for easy-level content. With more medium-difficulty sessions in the dataset, SMART-MLP can more easily identify patterns. Additionally, IRT models benefit from medium-difficulty items, as the variability in responses allows the model to better discriminate between skill levels, crucial for accurate estimates.
Table 3. Accuracy obtained for the proposed SMART-MLP model with 8-minutes time windows compared to the traditional IRT baseline, evaluated across different content difficulty levels in our contributed SMARTe-VR-db.
@@
5. Conclusion and Future Work
@@
In this work, we have: i) reviewed adaptive learning models and datasets; ii) introduced a VR platform for immersive education, specifically designed for adaptive learning, which monitors both biometric data and learning metadata; iii) released a public VR-based dataset for research in this area; iv) proposed five research challenges tied to this dataset; and v) designed two architectures for understanding detection to support dataset validation.
@@
We believe that VR will transform education by improving presence, interactivity, and biometric monitoring compared to traditional e-learning. SMARTe-VR provides a comprehensive VR framework equipped with tools such as Auto QA, textbook creation, and biometric monitoring, improving educator capabilities and enabling AI-driven advancements. The SMARTe-VR dataset and its associated challenges offer a valuable resource for advancing AL models in VR, supporting research in IRT, KT, and MMLA.
@@
SMART models offer an approach based on state-of-the-art IRT models, with a focus on understanding detection during learning. Using facial biometrics, they improve adaptability, enable real-time adjustments based on student states, and reduce the reliance on response data. These methods validated our data set with precision 85%, although more research is needed to ensure a wider applicability.
@@
In future work, we plan to expand the dataset with additional biometric modalities (Fierrez et al., 2018), more users, more questions per video, a broader range of lecture videos, and additional integrations of sensors and information sources (Acien et al., 2020b; Peña et al., 2023) to create a more robust resource. We also plan to monitor students in traditional e-learning settings under similar conditions to enable comparison between VR and traditional setups.
In parallel, we will address the security, privacy and vulnerability aspects of VR-based and avatar-mediated learning environments (Acien et al., 2020a; Pedrouzo-Rodriguez et al., 2025). As these systems become increasingly immersive and realistic, it will be critical to ensure robust identity management through biometric and behavioral analysis (Kobayashi et al., 2019; Jitpanyoyos et al., 2024). The growing scientific interest in the security of avatar-based systems (exemplified by the IJCB 2025 Special Session on Tackling Safety Issues in the Emerging Cybernetic Avatar Era) underscores the need for secure biometric authentication and comprehensive safety frameworks in immersive environments. Addressing these challenges will be an integral part of our future research to further strengthen the security and robustness of the SMARTe-VR platform.
@@
Furthermore, MR will be integrated into the platform to improve user experience. To address the challenges, we will share models with the community and refine the SMART-MLP and SMART-TCN models to detect understanding and include attention and fatigue detection. Ultimately, our goal is to incorporate these models into SMARTe-VR, creating a comprehensive adaptive learning experience that leverages the potential of VR-based education.
Finally, this work has been strongly focused on e-learning applications, but the developed technologies also find a good fit in many other application areas like in VR-based entertainment, simulation, and real-time operation of simulators and certain equipment (e.g., medical or industrial inspection).
Acknowledgements.
@@
This work was supported by the following projects: BBforTAI (PID2021-127641OB-I00 MICINN/FEDER), HumanCAIC (TED2021-131787B-I00 MICINN), BIO-PROCTORING (GNOSS Program, Agreement MINDEF-UAM-FUAM dated 29-03-2022), M2RAI (PID2024-160053OB-I00 MICIU/FEDER), and ELLIS Unit Madrid. A. Morales is also supported by the Madrid Government in the line of Excellence for University Teaching Staff (V PRICIT). This work was conducted during the doctoral research stay of Roberto Daza at the Nagao Laboratory, Nagoya University. The authors are particularly grateful to the experimenters from Nagoya University and Hiroshima City University for their significant contributions to data collection and experiment implementation. In addition, we extend our thanks to both universities for providing experimental equipment and technical support, which were indispensable for the successful completion of this research. Without their support, this study would not have been possible.